The endpoints of the diameter of a circle have coordinates R(-8,2) and S(10,-4). Find the coordinates of the center of the circle.
Answer by funmath(2933)

The endpoints of the diameter of a circle have coordinates R(-8,2) and S(10,-4). Find the coordinates of the center of the circle.
The center would be the midpoint of the diameter.
The midpoint formula is M=((x1+x2)/2,(y1+y2)/2)
(x1,y1)=(-8,2) and (x2,y2)=(10,-4)
M=((-8+10)/2,(2+(-4))/2_
M=(2/2,-2/2)
M=(1,-1)
The coordinates of the midpoint are (1,-1)
Plot the points and see if this seams reasonable.
Happy Calculating!!!!